OpenStudy (anonymous):

When simplifying the expression 2/x + 2/x-1 + 2/ x+2 . The denominator is x(x+1)(x–1), what is the numerator?

OpenStudy (anonymous):

The denominator is actually x(x-)(x+2). The numerator is 2( x(x-)(x+2))+2( x(x-)(x+2))+2( x(x-)(x+2)) = 6( x(x-)(x+2))

OpenStudy (anonymous):

6x2 + 4x – 4 6x2 + 9x + 4 6x2 + 8x + 4 6x2 + 9x – 4

OpenStudy (anonymous):

these are the answer choices

OpenStudy (anonymous):

I'm sorry, I missed a number for the numerator. Numerator = 6( x(x-1)(x+2))

OpenStudy (anonymous):

did you see the choices i sent ? is that one of the simplified ?

OpenStudy (anonymous):

Ok, given the function (2/x)+(2/(x-1))+(2/(x+2)), you can simplify it like this: (1) you name same denominator: (x(x-1)(x+2)) (2) factor the denominator out you'll get (x^3+x^2-2x). Now multiply that to all of the numerator. Final answer: (2/x)+(2/(x-1))+(2/(x+2)) can be simplified to (6x^3+6x^2-12x)/((x(x-1)(x+2))